# Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany

### v1

@@ III.5.2 Moving Averaged Data
Figures 15(b), 15(d), 16(b), 16(d) show a comparison of the moving average of the closed shutter data versus complex value of $C_{{\rm{open,}}j}$ normalized by the average value of $C[n]$ for each search. This is meant to convey what the phase evolution of an actual signal would look like compare to the data measured during the closed shutter periods. Here it is apparent that the values of $C_{{\rm{open,}}j}$ are well offset from origin clustered, for the most part, near 1+0i1+0\,i. The difference between the $C_{{\rm{open,}}j}$ data measured at the veto detector, which in general has much smaller deviations from 1+0i1+0\,i than the data measured at the science detector, shows the effect that the detuning of the HPL from the resonance of the RC has on the complex phase. This effect is discussed in more detail in Section IV.1.4. Again, the open shutter calibration will compensate for these effects.
III.5.3 Time Evolution of Backgrounds
@@
Figure 18: The mean open shutter signal (a) and closed shutter background (b) for $\rm S_{{}_{\perp}}$ and open shutter signal (c) and closed shutter background (d) for $\rm S_{{}_{\parallel}}$ is shown in blue as a function of the integration time. The darker grey lines show the 1 $\sigma$ uncertainty (between the 16th and 84th percentile) and the lighter grey lines show the 2 $\sigma$ uncertainty interval (between percentiles 2.5 and 97.5).
@@ IV Performance of the Optical System
Using the open shutter periods to calibrate the science data means that measuring the parameters of the optical system, such as the coupling of the HPL to the RC, is not strictly necessary. However, these measurements are important for verifying the performance of the optical system and also help identify the cause of changes in the calibration series. Calculating the important metrics to assess the performance of the optical system, such as the resonant enhancement of the RC and the total power overlap between the HPL and the RC, required measuring a number of other parameters. Some of these were dynamic parameters that were measured during ‘maintenance’ periods during the run, while other static parameters were measured before and/or after the run. This section discusses how these measurements were made and gives a summary of the performance of the optical system during the science runs.
@@ IV.1 Regeneration Cavity Parameters
The RC is a critical component of the optical system. Its resonant enhancement factor and the power overlap between its eigenmode and the HPL directly affect the sensitivity of the experiment. A more complete characterization of the RC can be found in Ref. [17]. This section will focus on how the RC performance was characterized during the first science campaign and the results of that effort. An explanation of the resonant enhancement factor, transmissivity, and reflectivity of the RC is given in Section IV.1.3, along with a discussion of the measurements of these parameters during the science runs. The FSR of the RC defines the frequencies used in the PLLs as well as the demodulation frequencies used in the heterodyne detection system. Changes in the FSR can therefore detune any regenerated signal from the resonance of the RC thus reducing the longitudinal field overlap of this light and the cavity eigenmode. The coupling of these changes to the magnitude and phase of the longitudinal overlap are discussed in Section IV.1.4, as are measurements of the RC FSR and longitudinal overlap during the science runs. The transmissivity and reflectivity of the RC as a function of the longitudinal overlap $\eta_{\hat{z}}$, are then discussed in Section IV.1.4.
@@ IV.1.1 Resonant Enhancement
Assuming a signal field that is perfectly on resonance and in the spatial mode of the RC, the resonant enhancement factor of the RC, $\beta_{\!{}_{\rm RC}}$, can be defined as the ratio of the power in the signal field that exits the cavity at the mirror coupled to the detection system, to the signal power at the same location if the cavity were not there at all. $\beta_{\!{}_{\rm RC}}$ therefore gives a measure of the amplification of the science signal. In this case, the BSM field reconverts to an electromagnetic field inside the cavity and is stored there while being reflected back and forth between the mirrors. During this time the field experiences the cavity gain as it constructively interferes with the electromagnetic field that is newly reconverted from the BSM field. In this way, the field is injected via the reconversion of the BSM field to an electromagnetic field, rather than transmitting through an input coupling mirror. The electromagnetic field that has built up in the cavity is then transmitted by RC1 and directed to the science detector.  [^28] The cavity will actually suppress signals at frequencies not within its linewidth, but this is not of a concern as the bandwidth of regenerated signal should be much smaller than this if the optical system is functioning as intended. Without considering the effects of the spatial or longitudinal field overlap, $\beta_{\!{}_{\rm RC}}$ can be expressed as [22]
@@
In this equation $T_{\!{}_{\rm RC1}}$ is the power transmissivity of RC1, while $A$ is the total attenuation of the cavity circulating power per round trip. This is due to the combination of mirror transmissivities and other optical losses due to effects including scattering from the mirror surface roughness, absorption in the mirror coatings, and clipping losses on the combined free apertures of the magnet string and reflective surfaces of the mirrors. It can also be seen that the resonant enhancement factor of the cavity is equivalent to its power buildup factor, that is the ratio of the circulating power to injected power on resonance (assuming the power is being injected at the mirror considered the output coupler for the resonant enhancement factor).
Assessing the total attenuation that the power circulating in the cavity experiences through one round trip ($A$) is critical to knowing not only the resonant enhancement factor during the run, but a number of other parameters as well. Fortunately $A$ can be measured via a combination the cavity FSR $f_{0}$ and storage time $\tau$. The relationship between these parameters is given by
@@
Details on the technique for measuring the storage time and the values for $T_{\!{}_{\rm RC1}}$ and $T_{\!{}_{\rm RC2}}$ can be found in the Refs. [17] and [29]. To measure the cavity FSR $f_{0}$, during the run, the laser AL was PDH locked to the cavity. Then the PLL3 frequency was tuned near 45 FSRs, until the power transmitted by the cavity was at a maximum. The value of the PLL3 frequency at that point was then divided by 45 to give the FSR of the RC. The uncertainty in this measurement was 1 Hz, leading to an uncertainty on the FSR of 0.02 Hz (see Table 1).
The measurements of $\beta_{\!{}_{\rm RC}}$ during the maintenance periods are shown in Figures 18(a) and 18(b) for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ respectively. For $\rm S_{{}_{\parallel}}$, the black vertical lines in the plot indicate times at which the RC eigenmode position on the mirrors was aligned to improve the storage time and in turn the resonant enhancement factor (no such realignments were required for $\rm S_{{}_{\perp}}$). For $\rm S_{{}_{\perp}}$ the mean value of $\beta_{\!{}_{\rm RC}}$ associated with the closed shutter periods was 7150 with a range from 6810 to 7120 and RMS fluctuations from its mean value of 90. The uncertainty on the measurements of $\beta_{\!{}_{\rm RC}}$ was 80 and was a combination of the uncertainty on the transmissivity of RC1 and the uncertainty on the measurements of the storage time. For the closed shutter periods of $\rm S_{{}_{\parallel}}$ the mean value of $\beta_{\!{}_{\rm RC}}$ was estimated to be 6870 with a range from 6640 to 6910 with an RMS of 70. These measurements had the same uncertainty as those performed during $\rm S_{{}_{\perp}}$. While a value below 5900 was measured during one of the periods of $\rm S_{{}_{\parallel}}$, and a value above 7200 was measured during one of the periods of $\rm S_{{}_{\perp}}$, there were no closed shutter periods during time periods associated with these measurements. For this reason this data point is not considered in the range of data mentioned above.
@@ IV.1.2 Transmissivity
Assuming 100% spatial overlap, the transmissivity of the cavity can be expressed as [27]
@@
In this equation it is clear that due to the common factors $T_{\rm RC,max}$ can be expressed as the product of the resonant enhancement and the transmissivity of the flat cavity mirror on the COB (again ignoring the spatial or longitudinal field overlap). It should also be noted that $T_{\rm RC,max}$ is independent of which side the input field is incident on.
Calculating the maximum transmissivity of the RC on resonance (assuming 100% spatial overlap) is relatively trivial with measurements of the transmissivities of RC1 and RC2, along with the storage time and FSR, as it is just the product
@@
Like the resonant enhancement factor, using the approximation above leads to an error in the measured transmissivity on the order of 0.01%, which is well below other sources of error in the measurement. For $\rm S_{{}_{\perp}}$, the average nominal (on resonance) transmissivity of the RC during the closed shutter periods was estimated to be 3.6% with a range from 3.5% to 3.6%. and RMS deviations from the mean of 0.05%. The uncertainty of $T_{\!{}_{\rm RC2}}$ was the primary contributor to the absolute uncertainty of $\pm 0.4\%$ on $T_{\rm RC,max}$. During the closed shutter periods of $\rm S_{{}_{\parallel}}$ the average nominal transmissivity of the RC was estimated to be 3.5%, with a range from 3.4% to 3.5%, an RMS of 0.04%, and an absolute uncertainty of $\pm 0.3\%$.
Measurements of the off-resonance transmissivity of the RC (again assuming 100% spatial overlap) are discussed in Section IV.1.4 as they involve the detuning of the HPL frequency from the RC resonance, $\Delta\nu$.
@@ IV.1.3 Reflectivity
Unlike the transmissivity, the reflectivity differs depending on the side of the cavity on which the laser is incident. Assuming that the input laser is perfectly mode matched to the cavity eigenmode, the reflectivities $R_{\rm RC,CH,min}$ for a laser incident on the flat cavity mirror in CH and $R_{\rm RC,NR,min}$ for a laser incident on the curved cavity mirror in NR are [27]:
@@
$R_{\rm RC,NR}$ $=\left|\frac{1}{R_{\!{}_{\rm RC1}}}\left(R_{\!{}_{\rm RC1}}-\frac{T_{\!{}_{\rm RC1}}}{1-\sqrt{1-A}\,e^{i2\pi\frac{\Delta\nu}{f_{0}}}}\right)\right|^{2}\simeq\left|\left(1-\frac{2T_{\!{}_{\rm RC1}}}{A}\frac{1}{1-i\frac{\Delta\nu}{\nu_{0}}}\right)\right|^{2},{\rm\hskip 8.5359ptand}$ (55)
$R_{\rm RC,CH}$ $=\left|\frac{1}{R_{\!{}_{\rm RC2}}}\left(R_{\!{}_{\rm RC2}}-\frac{T_{\!{}_{\rm RC2}}}{1-\sqrt{1-A}\,e^{i2\pi\frac{\Delta\nu}{f_{0}}}}\right)\right|^{2}\simeq\left|\left(1-\frac{2T_{\!{}_{\rm RC2}}}{A}\frac{1}{1-i\frac{\Delta\nu}{\nu_{0}}}\right)\right|^{2}.$ (56)
In both cases, the approximations assume $R_{\!{}_{\rm RC1}}$ and $R_{\!{}_{\rm RC2}}$ are one. Assuming that the HPL frequency satisfies the resonance condition of the cavity ($\Delta\nu\ll Af_{0}$), the minimum reflectivity of the RC assuming 100% spatial overlap can be approximated as
$R_{\rm RC,NR,min}$ $\simeq\left(1-\frac{2T_{\!{}_{\rm RC1}}}{A}\right)^{2},{\rm\hskip 8.5359ptand}$ (57)
$R_{\rm RC,CH,min}$ $\simeq\left(1-\frac{2T_{\!{}_{\rm RC2}}}{A}\right)^{2}.$ (58)
The reflectivity of the cavity from NR in terms of measured parameters is therefore given by
$R_{\rm RC,NR,min}\simeq\left(1-T_{\!{}_{\rm RC1}}f_{0}\tau\right)^{2},$ (59)
and from CH
$R_{\rm RC,CH,min}\simeq\left(1-T_{\!{}_{\rm RC2}}f_{0}\tau\right)^{2}.$ (60)
In this case, the approximation leads to a relative error of on the order of 0.01% and thus is not a significant source of error.
The off-resonant reflectivity assuming 100% spatial overlap is discussed in Section IV.1.4 as the detuning of the HPL frequency with respect to the resonance of the RC $\Delta\nu$, is also discussed there.
IV.1.4 Longitudinal Overlap

### v2

@@ III.5.2 Moving Averaged Data
Figures 15(b), 15(d), 16(b), 16(d) show a comparison of the moving average of the closed shutter data versus complex value of $C_{{\rm{open,}}j}$ normalized by the average value of $C[n]$ for each search. This is meant to convey what the phase evolution of an actual signal would look like compared to the data measured during the closed shutter periods. Here it is apparent that the values of $C_{{\rm{open,}}j}$ are well offset from origin clustered, for the most part, near 1+0i1+0\,i. The difference between the $C_{{\rm{open,}}j}$ data measured at the veto detector, which in general has much smaller deviations from 1+0i1+0\,i than the data measured at the science detector, shows the effect that the detuning of the HPL from the resonance of the RC has on the complex phase. This effect is discussed in more detail in Section IV.2.1. Again, the open shutter calibration will compensate for these effects.
@@
Figure 18: The mean open shutter signal (a) and closed shutter background (b) for $\rm S_{{}_{\perp}}$ and open shutter signal (c) and closed shutter background (d) for $\rm S_{{}_{\parallel}}$ is shown in blue as a function of the integration time. The black lines show the 1 $\sigma$ uncertainty (between the 16% and 84%) and the gray lines show the 2 $\sigma$ uncertainty interval (between 2.5% and 97.5%).
III.5.3 Time Evolution of Backgrounds
@@
III.6 Central Optical Bench Transmissivity
Precise measurements of combined transmissivity of the COB mirrors mounted on the optical axis of the system (see Figure 6),
$T_{{}_{\rm COB}}=T_{{}_{\rm LT1}}T_{{}_{\rm MZ3}}T_{{}_{\rm MZ4}}T_{{}_{\rm LT2}},$ (45)
are essential to calibrate the ratio of the closed to open shutter power in terms of the conversion rate between electromagnetic and BSM fields $\mathcal{P}_{\gamma\leftrightarrow\phi}$. This can seen in Equation 25 with the critical parameter being $T_{\!{}_{\rm COB}}T_{\!{}_{\rm RC2}}$. $T_{\!{}_{\rm COB}}$ is also an important to calculating the field overlap $\eta$ as Section IV.2 will explain.
Ex-situ measurements of these optics were performed after the completion of the first science campaign with the COB removed from the vacuum system. As the transmissivities of these optics were strongly dependent on the angle of incidence ($\sim$ 1.5 ppm per degree for the mirrors sealing the light-tight housings and $\sim$ 150 ppm per degree for the MZ mirrors) the flat cavity mirror of the RC was used as an alignment reference to ensure that the laser used in the setup passed through the optics at the same angle as the HPL light during the open shutter periods of the science runs. The measurement procedure involved shining a laser coupled to an optical fiber on the mirror being tested. The light reflected from the optic was then incident normal to the RC flat mirror and reflected back through the fiber. This ensured that the incident beam was aligned to the angle of the RC eigenmode (assuming no changes in the alignment of the mirror after the science campaign) to better than 1 mrad. The results are shown in Table 5. Similar results were expected for the pair of light tight housing mirrors and the pair of Mach-Zehnder mirrors, and the measurements of LT1 and LT2 both resulted in a transmissivity of $12.4\pm 0.5\rm\,ppm$. MZ3 was measured to have a transmissivity of $452\pm 10\rm\,ppm$ while the transmissivity of MZ4 was $275\pm 10\rm\,ppm$. While MZ3 and MZ4 were expected to have the same transmissivity as their dielectric coatings were applied during the same coating run, the discrepancy between them is believed to be related to differences in their angles of incidence with respect to the optical axis of the system. The uncertainty in these measurements is due to a combination of the uncertainty in the calibration of the power meters used in the setup, the presence of stray light at the power meters measuring the weak transmitted signals, and the uncertainty of the angle of incidence of the laser on the mirrors.
In-situ measurements of the on-axis COB optics were also performed in between $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. Here, the transmissivity of LT1 was observed via a measurement of the ratio of HPL power directly at the input window to the vacuum system in NL and then at the High-Power port of the CH vacuum tank (see Figure 6). This gave a value of $11\pm 1$ ppm. The transmissivity of LT2 was measured with a similar method, but instead measuring the RL power just before the Injection Port at the CH tank, also giving a value of $11\pm 1$ ppm. The combined transmissivity of MZ3 and MZ4 was measured using a power meter to measure the HPL power at the PLL2 port, and then at the Mach-Zehnder port by measuring the amplitude of the RL-HPL beatnote. The spatial overlap of the RL-HPL beatnote at optimal alignment could be directly measured at the PLL2 port because it was possible to measure the static power of both lasers there. This value could then be used to compensate for these effects in the measurement of the beatnote amplitude at the Mach-Zehnder port. [^28] It should be noted that to find the beatnote amplitude at each port for the ‘optimal’ alignment, RL was realigned to maximize the the beatnote amplitude in both cases. This was necessary to compensate for the alignment drift of the MZ on the COB. The combined transmissivity of the MZ mirrors found in this measurement was $0.137\pm 0.007$ ppm.
The result of the combined transmissivities of the MZ mirror from the ex-situ measurements gave $0.124\pm 0.005$ ppm. Thus, it is clear that the individual measurements of each of the on-axis COB optics agree between the two methods. Even more relevant is the total transmissivity $T_{{}_{\rm COB}}$, in which a value of $(1.9\pm 0.1)\times 10^{-17}$ was found for the ex-situ measurement with $(1.7\pm 0.2)\times 10^{-17}$ measured for the in-situ measurements. In short, these measurements agree within their uncertainty.
Finally, the most relevant parameter for the calibration of the results in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$ is the product of the transmissivity of COB optics and the cavity mirror RC2. For this, the ex-situ measurements gave $T_{{}_{\rm COB}}T_{{}_{\rm RC2}}=(9.7\pm 1.2)\times 10^{-23}$ while the in-situ results were $T_{{}_{\rm COB}}T_{{}_{\rm RC2}}=(8.5\pm 1.4)\times 10^{-23}$. Again, the measurements agree within their uncertainty. This gives us confidence in the accuracy of the overall calibration of the results. We consider the method for the ex-situ measurements to be more robust as we were able to directly access the optics. Therefore, the ex-situ measurements were used for the actual calibration of the results.
@@ IV Performance of the Optical System
Using the open shutter periods to calibrate the science data means that measuring the parameters of the optical system, such as the coupling of the HPL to the RC, is not strictly necessary. However, these measurements are important for verifying the performance of the optical system and also help identify the cause of changes in the calibration series. Calculating the important metrics to assess the performance of the optical system, such as the resonant enhancement of the RC and the total power overlap between the HPL and the RC, required measuring a number of other parameters. Some of these were dynamic parameters that were measured during maintenance periods during the run, while other static parameters were measured before and/or after the run. This section discusses how these measurements were made and gives a summary of the performance of the optical system during the science runs.
@@ IV.1 Regeneration Cavity Parameters
The RC is a critical component of the optical system. Its resonant enhancement factor and the power overlap between its eigenmode and the HPL directly affect the sensitivity of the experiment. A more complete characterization of the RC can be found in Ref. [17]. This section will focus on how the RC performance was characterized during the first science campaign and the results of that effort. Explanations of the resonant enhancement factor, transmissivity, and reflectivity of the RC are given in Sections IV.1.1, IV.1.2, and  IV.1.3 respectively, along with a discussion of the measurements of these parameters during the science runs.
@@ IV.1.1 Resonant Enhancement
The resonant enhancement factor of the RC, $\beta_{\!{}_{\rm RC}}$, can be defined as the ratio of the power in the signal field that exits the cavity at the mirror coupled to the detection system, to the signal power at the same location if the cavity were not there at all, for a signal field that is on resonance and in the spatial mode of the RC,. $\beta_{\!{}_{\rm RC}}$ therefore gives a measure of the amplification of the science signal. In this case, the BSM field reconverts to an electromagnetic field inside the cavity and is stored there while being reflected back and forth between the mirrors. During this time the field experiences the cavity gain as it constructively interferes with the electromagnetic field that is newly reconverted from the BSM field. In this way, the field is injected via the reconversion of the BSM field to an electromagnetic field, rather than transmitting through an input coupling mirror. The electromagnetic field that has built up in the cavity is then transmitted by RC1 and directed to the science detector. The resonant enhancement, $\beta_{\!{}_{\rm RC}}$, can be expressed as [22]
@@
In this equation, $T_{\!{}_{\rm RC1}}$ is the power transmissivity of RC1 while $A$ is the total round-trip attenuation of the power circulating in the cavity. This is due to the combination of mirror transmissivities and other optical losses including scattering from the mirror surface roughness, absorption in the mirror coatings, and clipping losses on the combined free apertures of the magnet string and reflective surfaces of the mirrors. The approximation assumes $A\ll 1$ and therefore $\sqrt{1-A}\simeq 1-A/2$, leading to an error in the resonant enhancement on the order of 0.01%.
Assessing the total attenuation $A$ is critical to knowing not only the resonant enhancement factor during the run, but a number of other parameters as well. Fortunately, $A$ can be measured via a combination the cavity FSR $f_{0}$ and storage time $\tau$. The relationship between these parameters is
@@
Details on the technique for measuring the storage time and the values for $T_{\!{}_{\rm RC1}}$ and $T_{\!{}_{\rm RC2}}$ can be found in the Refs. [17] and [29]. To measure the cavity FSR $f_{0}$ during the run, the laser AL was PDH locked to the cavity. Then the PLL3 frequency was tuned near 45 FSRs, until the power transmitted by the cavity was at a maximum. The value of the PLL3 frequency at that point was then divided by 45 to give the FSR of the RC. The uncertainty in this measurement was 1 Hz, leading to an uncertainty on the FSR of 0.02 Hz (see Table 1).
The measurements of $\beta_{\!{}_{\rm RC}}$ during the maintenance periods are shown in Figures 18(a) and 18(b) for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ respectively. For $\rm S_{{}_{\parallel}}$, the black vertical lines in the plot indicate times at which the RC eigenmode position on the mirrors was aligned to improve the storage time and in turn the resonant enhancement factor (no such realignments were required for $\rm S_{{}_{\perp}}$). For $\rm S_{{}_{\perp}}$ the mean value of $\beta_{\!{}_{\rm RC}}$ associated with the closed shutter periods was 7150 with a range from 6810 to 7120 and RMS fluctuations from its mean value of 90. The uncertainty on the measurements of $\beta_{\!{}_{\rm RC}}$ was 80, mostly due to the uncertainty on the transmissivity of RC1 and the uncertainty on the measurements of the storage time. For the closed shutter periods of $\rm S_{{}_{\parallel}}$ the mean value of $\beta_{\!{}_{\rm RC}}$ was estimated to be 6870 with a range from 6640 to 6910 with an RMS of 70. These measurements had the same uncertainty as those performed during $\rm S_{{}_{\perp}}$. While a value below 5900 was measured during one of the periods of $\rm S_{{}_{\parallel}}$, and a value above 7200 was measured during one of the periods of $\rm S_{{}_{\perp}}$, there were no closed shutter periods during time periods associated with these measurements. For this reason this data point is not considered in the range of data mentioned above.
@@ IV.1.2 Transmissivity
The transmissivity of the cavity for a laser occupying its spatial eigenmode (the spatial coupling will be examined later) can be expressed as [27]
@@
In this equation it is clear that due to the common factors, $T_{\rm RC,max}$ can be expressed as the product of the resonant enhancement and the transmissivity of the flat cavity mirror on the COB (again ignoring the spatial or longitudinal field overlap). It should also be noted that $T_{\rm RC,max}$ is independent of which side the input field is incident on. The maximum transmissivity of the RC on resonance can therefore be expressed as the product
@@
For $\rm S_{{}_{\perp}}$, during the closed shutter periods $T_{\rm RC,max}$ was estimated to have an average value of 3.6% with a range from 3.5% to 3.6% and RMS deviations from the mean of 0.05%. The uncertainty of $T_{\!{}_{\rm RC2}}$ was the primary contributor to the absolute uncertainty of $\pm 0.4\%$ on $T_{\rm RC,max}$. During the closed shutter periods of $\rm S_{{}_{\parallel}}$ the average of $T_{\rm RC,max}$ was estimated to be 3.5%, with a range from 3.4% to 3.5%, an RMS of 0.04%, and an absolute uncertainty of $\pm 0.3\%$.
As Equation 53 shows, detuning the incident laser frequency from the cavity resonance will reduce its transmissivity. Measurements of the detuning $\Delta\nu$ between the HPL frequency and resonance of the RC are discussed in Section IV.2.1 and were used to infer the transmissivity of the RC during the first science campaign. During the closed-shutter periods of $\rm S_{{}_{\perp}}$, $T_{\rm RC}$ had a mean value of 3.3%, with RMS fluctuations about the mean of 0.2%, a maximum value of 3.6%, and a minimum value of 2.6%. The uncertainty of these measurements was 0.3%. For $\rm S_{{}_{\perp}}$, the closed-shutter periods gave a mean value of 3.0% for $T_{\rm RC}$, with RMS fluctuations of 0.5%, a maximum of 3.5%, and a minimum of 1.4%. The uncertainty was the same as the measurements during $\rm S_{{}_{\perp}}$ at 0.3%. In both runs the uncertainty was almost entirely driven by the uncertainty in the measurement of the transmissivity of the flat RC mirror mounted to the COB, RC2.
@@ IV.1.3 Reflectivity
The reflectivity of the RC is used in the calibration of several critical parameters used to evaulate the performance of the optical system. Unlike the transmissivity, the reflectivity differs depending on the side of the cavity on which the laser is incident. The reflectivity experienced by the HPL when incident on the flat RC mirror in the CH cleanroom ($\rm RC2$) will be referred to as $R_{\rm RC,CH}$ while the reflectivity for lasers incident on the curved cavity mirror in the NR cleanroom ($\rm RC1$) will be referred to as $R_{\rm RC,NR}$. These quantities can be expressed as [27]
@@
$R_{\rm RC,CH}$ $=\left|\frac{1}{R_{\!{}_{\rm RC2}}}\left(R_{\!{}_{\rm RC2}}-\frac{T_{\!{}_{\rm RC2}}}{1-\sqrt{1-A}\,e^{i2\pi\frac{\Delta\nu}{f_{0}}}}\right)\right|^{2}\simeq\left|\left(1-\frac{2T_{\!{}_{\rm RC2}}}{A}\frac{1}{1-i\frac{\Delta\nu}{\nu_{0}}}\right)\right|^{2},{\rm\hskip 8.5359ptand}$ (56)
$R_{\rm RC,NR}$ $=\left|\frac{1}{R_{\!{}_{\rm RC1}}}\left(R_{\!{}_{\rm RC1}}-\frac{T_{\!{}_{\rm RC1}}}{1-\sqrt{1-A}\,e^{i2\pi\frac{\Delta\nu}{f_{0}}}}\right)\right|^{2}\simeq\left|\left(1-\frac{2T_{\!{}_{\rm RC1}}}{A}\frac{1}{1-i\frac{\Delta\nu}{\nu_{0}}}\right)\right|^{2}.$ (57)
In both cases, the approximations assume the reflectivity of $\rm RC1$ and $\rm RC2$ are one. This can be expressed in terms of measured parameters as
$R_{{}_{\rm RC,CH}}$ $\simeq\left|1-\frac{T_{{}_{\rm RC2}}f_{0}\tau}{1-i2\pi\tau\Delta\nu}\right|^{2},{\rm\hskip 8.5359ptand}$ (58)
$R_{{}_{\rm RC,NR}}$ $\simeq\left|1-\frac{T_{{}_{\rm RC1}}f_{0}\tau}{1-i2\pi\tau\Delta\nu}\right|^{2}.$ (59)
In Sections IV.2.4 and IV.2.5 $R_{{}_{\rm RC,CH}}$ is used to calibrate the spatial overlap and field overlap measured at the veto detection system respectively. During the closed shutter sections of $\rm S_{{}_{\perp}}$ average value of $R_{{}_{\rm RC,CH}}$ was 92%, with a range from 92% to 94%, and RMS fluctuations from the mean of 0.6%. For $\rm S_{{}_{\parallel}}$ the RC reflectivity gave an average value of 93%, with values ranging between 92% to 97%, and an RMS of 1%. The uncertainty of these measurements was 1%, largely driven by the uncertainty in measurements of the transmissivity of RC2.
The reflectivity of the RC for the AL is used in Section  to calculate its spatial overlap with the cavity eigenmode. Since AL is injected to the RC via $\rm RC1$ and is frequency stabilized directly to a cavity resonance, Equation 59 for $R_{{}_{\rm RC,NR}}$ can be used with $\Delta\nu=0$. In this case the reflectivity of the cavity had an average value of 2.6% during the closed shutter periods of $\rm S_{{}_{\perp}}$ with a range from 2.4% to 3.1% and an RMS of 0.2%. During $\rm S_{{}_{\parallel}}$ the average reflectivity was 3.0% while the range was from 2.8% to 3.4%. The RMS of the reflectivity during these measurements was 0.2%. For both runs the absolute uncertainty in the reflectivity was 1.3% and was due to the uncertainty in the measurements of the cavity storage time and the transmissivity of $\rm RC1$.
IV.2 Field overlap
The field overlap between the HPL and the RC, $\eta$, is a critical metric for evaluating the performance of the optical system. As explained in Section I.1, it can be expressed as the product of spectral ($\eta_{\hat{z}}$), polarization ($\eta_{\rm pol}$), and spatial ($\eta_{xy}$) components.
The spectral component, also called the longitudinal overlap, is discussed in Section IV.2.1 and can be calculated exclusively from measurements of the RC storage time and the frequency detuning of HPL from its resonance. The polarization overlap, discussed in Section IV.2.2, is calculated from the orientation of polarization state of the HPL with respect to the magnetic field.
The spatial overlap is measured by calibrating the results of the open shutter periods in terms of the HPL power leaving the cavity $P_{{\rm open}}$, and dividing by the expected HPL power at this point for a laser occupying the spatial eigenmode of the RC. At the science detector, this can be expressed as
$\eta_{{xy}}^{2}=\frac{P_{{\rm open}}}{T_{\!{}_{\rm COB}}T_{\!{}_{\rm RC}}P_{\rm i}}.$ (60)
The expected power is therefore the product of the combined transmissivities of the COB optics $T_{\!{}_{\rm COB}}$, the transmissivity of the RC, $T_{\!{}_{\rm RC}}$, and the HPL power traveling though the production magnet string $P_{\rm i}$.
The veto detector can also be used to measure the field overlap between the HPL and the RC. There are even some advantages to using the veto detector as the HPL does not need to pass through the cavity to reach it resulting in a HPL power is higher at the veto PD. Therefore, the following sections include a discussion of the field overlap measured at both the science and veto detectors. The process of measuring the ratio $P_{{\rm open}}/P_{{\rm i}}$ is explained in Section IV.2.3. Projections of the spatial overlap during the first science campaign are discussed in Section IV.2.4, while the total field overlap $\eta$ is discussed in Section IV.2.5. The methods used to optimize the spatial overlap are described in Section IV.2.6.
IV.2.1 Longitudinal Overlap